# On a discrete version of the position-momentum commutation relation

### v4

@@ Abstract
The usual position-momentum commutation relation plays a fundamental role in the mathematical description of continuous-variable quantum systems. In the case of a qudit described by a Hilbert space of a high enough dimension, there exists a discrete version of the position-momentum commutation relation approximately satisfied by a large part of the pure quantum states. Our purpose is to explore in more details the set of these states. We show that it contains a family of discrete-variable Gaussian states depending on a continuous parameter and certain discrete coherent states. It also contains various discrete-variable versions of the Hermite-Gauss states, defined either as eigenstates of certain discrete versions of the harmonic oscillator Hamiltonian or generated by using a discrete version of the creation or annihilation operator. As a direct consequence, a discrete version of the incertitude relation is satisfied by the considered quantum states.
preprint: APS/123-QED
@@ I Introduction
The quantum systems with finite-dimensional Hilbert space [1, 2, 3, 4, 5] play an important role in the investigation of quantum mechanics foundations, in Quantum Information, and in other applications. The mathematical description of these discrete-variable systems is obtained by following, as much as possible, the analogy with the continuous-variable quantum systems. In certain cases, the discrete-variable versions have properties very similar to those concerning the continuous systems, but this does not happen in all the cases.
In quantum mechanics, a pair $A$ and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In [6], the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in [6].
Our approach is different. We show that a discrete version of the position-momentum commutation relation $\bar{q},\bar{p}]\!=\!\mathrm{i}\hbar$ is approximately satisfied on a proper subspace if the dimension of the Hilbert space is high enough. Among the elements of such a subspace there are some of the most important states used in the mathematical description of qudits [7, 8, 9, 10].
The presented results may be used as a starting point of new mathematical formalisms involving only certain particular states of a qudit, by following, for example, the analogy with the theory of continuous-variable Gaussian states used in Quantum Information [11, 12, 13].
@@ II Quantum systems with finite-dimensional Hilbert space
The discrete $\delta$ functions $\delta_{-s}$, $\delta_{-s+1}$, … , $\delta_{s}$, where
@@
III Discrete Fourier transform
The usual Fourier transform
$\begin{array}[]{l}\hat{F}\!:\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{F}\Psi,\\[5.69054pt] (\hat{F}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{h}pq]\,\Psi(q)\,dq,\end{array}$ (8)
is a unitary transform. Its inverse is the adjoint transform
$\begin{array}[]{l}\hat{F}^{\dagger}\!:\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{F}^{\dagger}\Psi,\\[5.69054pt] (\hat{F}^{\dagger}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{h}pq]\,\Psi(q)\,dq,\end{array}$ (9)
The corresponding discrete-variable version [2, 7],
$\begin{array}[]{l}\hat{\mathfrak{F}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{F}}\psi,\\ (\hat{\mathfrak{F}}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{d}kn]\,\psi(n),\end{array}$ (10)
is also a unitary transform. Its inverse is the adjoint
$\begin{array}[]{l}\hat{\mathfrak{F}}^{\dagger}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{F}}^{\dagger}\psi,\\ (\hat{\mathfrak{F}}^{\dagger}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}kn]\,\psi(n).\end{array}$ (11)
IV Discrete version of position and momentum operators
The position operator
$\begin{array}[]{l}\hat{q}\!:\!D_{q}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{q}\,\Psi,\\[2.84526pt] (\hat{q}\,\Psi)(q)\!=\!q\,\Psi(q),\end{array}$ (12)
@@ IV Discrete version of position and momentum operators
admits the discrete-variable version [1, 2]
$\hat{\mathfrak{q}}:\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{q}}\psi,\quad(\hat{\mathfrak{q}}\psi)(n)\!=\!n\,\psi(n).$ (14)
The basis $\{|n\rangle\}_{n=-s}^{s}$ is an eigenbasis of $\hat{\mathfrak{q}}$, and $\hat{\mathfrak{q}}$ admits the spectral resolution
$\hat{\mathfrak{q}}\!=\!\sum\limits_{n=-s}^{s}n|n\rangle\langle n|.$ (15)
The momentum operator
$\begin{array}[]{l}\hat{p}\!:\!D_{p}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{p}\,\Psi,\\[2.84526pt] \hat{p}\,\Psi\!=\!-\mathrm{i}\hbar\frac{\mathrm{d}}{\mathrm{d}q}\Psi,\end{array}$ (16)
@@
$\hat{p}\!=\!\hat{F}^{\dagger}\hat{q}\hat{F}$. (18)
The discrete-variable version of momentum is defined as [1, 2]
$\hat{\mathfrak{p}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\quad\hat{\mathfrak{p}}\!=\!\hat{\mathfrak{F}}^{\dagger}\hat{\mathfrak{q}}\hat{\mathfrak{F}}$. (19)
@@
$\begin{array}[]{l}(\hat{\mathfrak{p}}\psi)(n)\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}\sum\limits_{m=-s}^{s}k\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k(n-m)]\,\psi(m),\end{array}$ (20)
and $\hat{\mathfrak{p}}$ admits the spectral resolution
$\hat{\mathfrak{p}}\!=\!\sum\limits_{n=-s}^{s}n\,\hat{\mathfrak{F}}^{\dagger}|n\rangle\langle n|\hat{\mathfrak{F}}$, (21)
that is $\{\hat{\mathfrak{F}}^{\dagger}|n\rangle\}_{n=-s}^{s}$ is an eigenbasis of $\hat{\mathfrak{p}}$.
V Discrete version of the position-momentum commutation relation
Table 1: Eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ in two cases.

### v5

@@ Abstract
The usual position-momentum commutation relation plays a fundamental role in the mathematical description of continuous-variable quantum systems. In the case of a qudit described by a Hilbert space of a high enough dimension, there exists a discrete version of the position-momentum commutation relation satisfied with approximation of a large part of the pure quantum states. Our purpose is to explore in more details the set of these states. We show that it contains a family of discrete-variable Gaussian states depending on a continuous parameter and certain discrete coherent states. It also contains various discrete-variable versions of the Hermite-Gauss states, defined either as eigenstates of certain discrete versions of the harmonic oscillator Hamiltonian or generated by using a discrete version of the creation or annihilation operator. The presented results suggest a possible way to experimentally obtain some of the investigated states.
preprint: AIP/123-QED
@@ I Introduction
The quantum systems with finite-dimensional Hilbert space Schwinger60 ; Vourdas04 ; Cotfas11 ; Cotfas23 ; Cotfas24 play an important role in the investigation of quantum mechanics foundations, in Quantum Information, and in other applications. The mathematical description of these discrete-variable systems is obtained by following, as much as possible, the analogy with the continuous-variable quantum systems. In certain cases, the discrete-variable versions have properties very similar to those concerning the continuous systems, but this does not happen in all the cases.
In quantum mechanics, a pair $A$ and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In Farrales25 , the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in Farrales25 .
Our approach is different. We show that a discrete version of the position-momentum commutation relation $\bar{q},\bar{p}]\!=\!\mathrm{i}\hbar$ is satisfied with an approximation in a proper subspace if the dimension of the Hilbert space is high enough. Among the elements of such a subspace there are some of the most important states used in the mathematical description of qudits Mehta87 ; Barker ; Nikiforov ; Vilenkin .
The presented results may be used as a starting point of new mathematical formalisms involving only certain particular states of a qudit, by following, for example, the analogy with the theory of continuous-variable Gaussian states used in Quantum Information Ferraro05 ; Wang07 ; Adesso14 .
@@ II Quantum systems with finite-dimensional Hilbert space
The set $\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ contains one and only one representative of any class of remainders modulo $d$. Therefore, we can identify it with $\mathbb{Z}_{d}$. The discrete $\delta$ functions $\delta_{-s}$, $\delta_{-s+1}$, … , $\delta_{s}$, where $\delta_{m}\!:\!\{-\!s,\!-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{C},$
@@
II.1 Discrete Fourier transform
The usual Fourier transform $\Psi\!\mapsto F\Psi,$
$\begin{array}[]{l}(F\Psi)(p)\!=\!\frac{1}{\sqrt{2\pi\hbar}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{-\frac{\mathrm{i}}{\hbar}pq}\,\Psi(q)\,dq\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}pq}\,\Psi(q)\,dq,\end{array}$ (8)
is a unitary transform. Its inverse is the adjoint transform $\Psi\!\mapsto F^{\dagger}\Psi,$
$\begin{array}[]{l}(F^{\dagger}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}pq}\,\Psi(q)\,dq,\end{array}$ (9)
The corresponding discrete-variable version $\mathbf{F}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\mathbf{F}\psi,$
$(\mathbf{F}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,\psi(n),$ (10)
is also a unitary transform Vourdas04 ; Mehta87 . Its inverse is the adjoint $\mathbf{F}^{\dagger}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\mathbf{F}^{\dagger}\psi,$
$(\mathbf{F}^{\dagger}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}\,\psi(n).$ (11)
Particularly, $\langle n|\mathbf{F}|k\rangle\!=\!(\mathbf{F}\delta_{k})(n)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}$, $\langle n|\mathbf{F}^{\dagger}|k\rangle\!=\!(\mathbf{F}^{\dagger}\delta_{k})(n)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}$, and
$\begin{array}[]{l}\mathbf{F}\!=\!\mathbb{I}\mathbf{F}\mathbb{I}\!=\!\sum_{n,k=-s}^{s}|n\rangle\langle n|\mathbf{F}|k\rangle\langle k|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,|n\rangle\langle k|,\\[2.84526pt] \mathbf{F}^{\dagger}\!=\!\mathbb{I}\mathbf{F}^{\dagger}\mathbb{I}\!=\!\sum_{n,k=-s}^{s}|n\rangle\langle n|\mathbf{F}^{\dagger}|k\rangle\langle k|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n,k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}\,|n\rangle\langle k|.\end{array}$ (12)
II.2 Discrete version of the position and momentum operators
The position operator $\hat{q}\!:\!D_{q}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{q}\,\Psi,$
$\begin{array}[]{l}(\hat{q}\,\Psi)(q)\!=\!q\,\Psi(q),\end{array}$ (13)
@@ II.2 Discrete version of the position and momentum operators
admits Schwinger60 ; Vourdas04 the discrete-variable version $\hat{\mathbf{q}}:\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathbf{q}}\psi,$
$(\hat{\mathbf{q}}\psi)(n)\!=\!n\,\psi(n).$ (15)
The basis $\{|n\rangle\}_{n=-s}^{s}$ is an eigenbasis for $\hat{\mathbf{q}}$, and $\hat{\mathbf{q}}$ admits the spectral decomposition
$\hat{\mathbf{q}}\!=\!\sum\limits_{n=-s}^{s}n|n\rangle\langle n|.$ (16)
The momentum operator $\hat{p}\!:\!D_{p}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{p}\,\Psi,$
$\begin{array}[]{l}\hat{p}\,\Psi\!=\!-\mathrm{i}\hbar\frac{\mathrm{d}}{\mathrm{d}q}\Psi,\end{array}$ (17)
@@
$\hat{p}\!=\!F^{\dagger}\hat{q}F.$ (19)
The discrete-variable version Schwinger60 ; Vourdas04 of momentum is defined as $\hat{\mathbf{p}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathbf{p}}\!=\!\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F}.$ (20)
@@
$\begin{array}[]{l}(\hat{\mathbf{p}}\psi)(n)\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}\sum\limits_{m=-s}^{s}k\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(n-m)}\,\psi(m),\end{array}$ (21)
and $\hat{\mathbf{p}}$ admits the spectral decomposition
$\hat{\mathbf{p}}\!=\!\sum\limits_{n=-s}^{s}n\,\mathbf{F}^{\dagger}|n\rangle\langle n|\mathbf{F},$ (22)
that is $\{\mathbf{F}^{\dagger}|n\rangle\}_{n=-s}^{s}$ is an eigenbasis for $\hat{\mathbf{p}}$.
II.3 A $d$-dimensional version of Pauli operators
The two-dimensional Pauli operator $Z:\mathbb{C}^{2}\!\longrightarrow\!\mathbb{C}^{2},$
$\begin{array}[]{l}Z|0\rangle\!=\!|0\rangle,\\ Z|1\rangle\!=\!-|1\rangle,\end{array},$ (23)
admits Vourdas04 the general version $\mathbf{Z}\!:\!\mathcal{H}\!\longrightarrow\!\mathcal{H},$
$\mathbf{Z}|n\rangle\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle$ (24)
satisfying the relations
@@ II.3 A $d$-dimensional version of Pauli operators
$\begin{array}[]{l}\mathbf{Z}\!=\!\mathbf{Z}\mathbb{I}\!=\!\mathbf{Z}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathbf{Z}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|,\\[5.69054pt] (\mathbf{Z}\psi)(n)\!=\!\langle n|\mathbf{Z}|\psi\rangle\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m}\langle n|m\rangle\langle m|\psi\rangle\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}\psi(n),\\[5.69054pt] \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\mathbb{I}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|\!=\!\mathbf{Z}.\end{array}$ (25)
The two-dimensional Pauli operator $X:\mathbb{C}^{2}\!\longrightarrow\!\mathbb{C}^{2},$
$\begin{array}[]{l}X|0\rangle\!=\!|1\rangle,\\ X|1\rangle\!=\!|0\rangle,\end{array}$ (26)
admits Vourdas04 the general version $\mathbf{X}\!:\!\mathcal{H}\!\longrightarrow\!\mathcal{H},$
$\mathbf{X}|n\rangle\!=\!|n\!+\!1\rangle$ (27)
where the addition is modulo $d$, and we choose $n\!+\!1\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$. It has the properties
$\begin{array}[]{l}\mathbf{X}\!=\!\mathbf{X}\mathbb{I}\!=\!\mathbf{X}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathbf{X}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}|n\!+\!1\rangle\langle n|,\\[5.69054pt] (\mathbf{X}\psi)(n)\!=\!\langle n|\mathbf{X}|\psi\rangle\!=\!\sum\limits_{m=-s}^{s}\langle n|m\!+\!1\rangle\langle m|\psi\rangle\!=\!\psi(n\!-\!1),\\[5.69054pt] \mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{p}}}\!=\!\mathbf{F}^{\dagger}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\mathbf{F}\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathbf{F}^{\dagger}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,|k\rangle\langle n|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathbf{F}^{\dagger}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k}|k\rangle\langle n|\\[2.84526pt] \qquad\ \ \ =\!\mbox{\small$\frac{1}{\sqrt{d}}$}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k(n+1)}\,\mathbf{F}^{\dagger}|k\rangle\langle n|\!=\!\mbox{\small$\frac{1}{d}$}\sum\limits_{n,k,m=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k(n+1)}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,|m\rangle\langle n|\\[2.84526pt] \qquad\ \ \ =\!\sum\limits_{n,m=-s}^{s}\mbox{\small$\frac{1}{d}$}\sum\limits_{k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(m-n-1)}\,|m\rangle\langle n|\!=\!\sum\limits_{n,m=-s}^{s}\delta_{m,n+1}\,|m\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}|n\!+\!1\rangle\langle n|\!=\!\mathbf{X}.\end{array}$ (28)
The operators $\mathbf{Z},\ \mathbf{X}:\!\mathcal{H}\!\longrightarrow\!\mathcal{H}$ are unitary transformations:
$\begin{array}[]{l}\langle\mathbf{Z}\varphi,\mathbf{Z}\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\mathbf{Z}\varphi(n)}\ \mathbf{Z}\psi(n)\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n)}\,\psi(n)\!=\!\langle\varphi,\psi\rangle,\\[2.84526pt] \langle\mathbf{X}\varphi,\mathbf{X}\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\mathbf{X}\varphi(n)}\ \mathbf{X}\psi(n)\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n\!-\!1)}\,\psi(n\!-\!1)\!=\!\langle\varphi,\psi\rangle.\end{array}$ (29)
II.4 A discrete version of the displacement operators
From
$\begin{array}[]{l}\mathbf{Z}\mathbf{X}\!=\!\sum\limits_{n,m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|m\!+\!1\rangle\langle m|\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}(m+1)}|m\!+\!1\rangle\langle m|,\\[2.84526pt] \mathbf{X}\mathbf{Z}\!=\!\sum\limits_{n,m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|m\!+\!1\rangle\langle m|n\rangle\langle n|\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m}|m\!+\!1\rangle\langle m|\end{array}$ (30)
it follows that
$\mathbf{Z}\mathbf{X}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}}\,\mathbf{X}\mathbf{Z},$ (31)
and the relations
[…]
In the continuous-variable case, the displacement operators are defined as
$\mathbf{D}(q,p)\!=\!\mathrm{e}^{-\frac{\mathrm{i}}{2\hbar}qp}\,\mathrm{e}^{\frac{\mathrm{i}}{\hbar}p\hat{q}}\,\mathrm{e}^{-\frac{\mathrm{i}}{\hbar}q\hat{p}}\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{h}qp}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}p\hat{q}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}q\hat{p}}$. (33)
By following the analogy with the continuous case, the discrete displacement operators are defined as $D(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\mathbf{D}(n,k)\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathbf{Z}^{k}\,\mathbf{X}^{n}\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k\hat{\mathbf{q}}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}n\hat{\mathbf{p}}}$. (34)
They satisfy the relations Vourdas04
$\begin{array}[]{l}\mathbf{D}(n,k)\psi(m)=\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,\psi(m\!-\!n),\\[5.69054pt] \mathbf{D}(n,k)\,\mathbf{D}(m,\ell)\!=\!\mathrm{e}^{\frac{\pi\mathrm{i}}{d}(km-n\ell)}\ \mathbf{D}(n\!+\!m,k\!+\!\ell)\\[5.69054pt] \mathbf{D}(n,k)\,\mathbf{D}(m,\ell)\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}(km-n\ell)}\ \mathbf{D}(m,\ell)\ \mathbf{D}(n,k)\\[5.69054pt] \mathbf{D}^{\dagger}(n,k)\!=\!\mathbf{D}(-n,-k)\!=\!\mathbf{D}^{-1}(n,k).\end{array}$ (35)
III Discrete version of the position-momentum commutation relation
Table 1: Eigenvalues of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ in two cases.